The equation of the chord of the circle x2 + y2 – 4x = 0 whose mid point is (1, 0) is
(A) y = 2 (B) y = 1 (C) x = 2 (D) x = 1

Grade:

2 Answers

Arun
3 years ago
Dear student
 
Use T = S1
 
x – 2 (x + 1) = 1- 4
 
 – x – 2 = – 3
 
x = 1
 
Hence option D is correct
 
Thanks and regards
Vikas TU
3 years ago
The equation of the chord is
x×1+0×y−2(x+1)=1+0−4 [ Using T=S1]
⇒−x−2=−3
⇒x=1
…...............................................